A spring having constant k is cut into n equal parts.All parts are connected into in series and then parallel.What is series to parallel combination ratio??

Grade:12

1 Answers

Vikas TU
7 years ago
in parallel
Keq.p = k + k + k + ….....+ nk = nk
in series
1/Keq.s = 1/k +1/k +1/k + …...+ n/k =n/k
Keq.s = k/n
ratio => nk*k/n =k*k = k^2